# Sigma[Uncaptioned image]: Siamese Mamba Network for Multi-Modal Semantic Segmentation

### v2

@@ 1 Introduction
Previous approaches in multi-modal semantic segmentation rely on Convolutional Neural Networks (CNN) or Vision Transformers (ViT). While CNN-based approaches [12, 24] are known for their scalability and linear complexity, they suffer from the small receptive field limited by the kernel size, leading to the local reductive bias. Besides, CNNs utilize a weight-sharing mechanism across different parts of the input, limiting its flexibility in adapting to unseen or low-quality images. In contrast, ViT-based methods [45, 25, 1] offer enhanced visual modeling by leveraging global receptive fields and dynamic weights. However, their self-attention mechanism leads to quadratic complexity in terms of input sizes [14], raising efficiency concerns. Attempts to improve this efficiency by reducing the dimensions or strides of the processing windows compromise the extent of the receptive fields [57].
@@
Our Sigma integrates a Siamese encoder for modality-specific feature extraction, fusion modules for aggregating information from different modalities, and a decoder for enhancing both spatial and channel-specific information. The encoder backbone utilizes cascaded Visual State Space Blocks (VSSB) with downsampling to extract multi-scale information from various modalities. Subsequently, the extracted features are inputted to a fusion module at each level, where multi-modal features initially interact through a Cross Mamba Block (CroMB) to augment cross-modal information. Then, the enhanced features are processed by a Concat Mamba Block (ConMB), which employs an attention mechanism to select pertinent information from each modality. Finally, the fused features are sent to multi-level Channel-Aware Visual State Space Blocks (CAVSSB) to effectively capture multi-scale long-range information.
@@ 3.4 Fusion Module
^$\hat{\mathcal{F}}_{\mathsf{rgb}}^{k},\;\hat{\mathcal{F}}_{\mathsf{x}}^{k}=\texttt{CroMB}(\mathcal{F}_{\mathsf{rgb}}^{k},\;\mathcal{F}_{\mathsf{x}}^{k}),$ (6)
@@
Here, ^$\hat{\mathcal{F}}_{\mathsf{rgb}}^{k},\hat{\mathcal{F}}_{\mathsf{x}}^{k},\text{and }\mathcal{F}_{\mathsf{fuse}}^{k}$ remains the original dimension as $\mathbb{R}^{H_{k}\times W_{k}\times C_{k}}$. Details of CroMB and ConMB are as follows.
@@
CroMB. As demonstrated in the upper part of Fig. 4, two input features are first processed by linear layers and depth-wise convolutions, then sent to the Cross Selective Scan (Cross SS) module. According to the selection mechanism of Mamba mentioned in Sec. 3.1, the system matrices $B,\;C$ and $\mathrm{\Delta}$ are generated to enable the context-aware ability of the model. Here, linear projection layers are utilized to generate the matrices. According to Eq. 4, matrix $C$ is used to decode the information from the hidden state $h_{k}$ to obtain the output $y_{k}$. Inspired by the cross-attention mechanism [3], we facilitate information exchange among multiple selective scan modules. In particular, the process can be represented as:
@@
Specifically, the outputs ^$\hat{\mathcal{F}}_{\mathsf{rgb}}^{k},\hat{\mathcal{F}}_{\mathsf{x}}^{k}\in\mathbb{R}^{H_{k}\times W_{k}\times C_{k}}$ from CroMB are first processed by linear and depth-wise convolution layers, then sent to the Concat Selective Scan (Concat SS) module. Within the Concat SS module, the two features are first flattened and then concatenated in the sequence length dimension. This provides a sequence $\mathcal{S}_{\mathsf{Concat}}^{k}$ of shape $\mathbb{R}^{(2\times H_{k}\times W_{k})\times C_{k}}$. Besides, to comprehensively capture information from two modalities, we inversely scan the concatenated sequence $\mathcal{S}_{\mathsf{Concat}}^{k}$ to obtain an additional sequence $\mathcal{S}_{\mathsf{Inverse}}^{k}\in\mathbb{R}^{(2\times H_{k}\times W_{k})\times C_{k}}$. Afterwards, each sequence is processed to capture long-range dependencies, obtaining $\hat{\mathcal{S}}_{\mathsf{Concat}}^{k}$ and $\hat{\mathcal{S}}_{\mathsf{Inverse}}^{k}$. Then the inversed sequence output is flipped back and added with the processed sequence. The summed sequence is separated to recover two outputs. This process can be represented as:
~$\widetilde{\mathcal{F}}_{\mathsf{rgb}}^{k}=\texttt{DWConv}(\texttt{Linear}(\hat{\mathcal{F}}_{\mathsf{rgb}}^{k})),$ (13)
~$\widetilde{\mathcal{F}}_{\mathsf{x}}^{k}=\texttt{DWConv}(\texttt{Linear}(\hat{\mathcal{F}}_{\mathsf{x}}^{k})),$ (14)
@@
¯$\overline{\mathcal{F}}_{\mathsf{RGB}}^{k},\overline{\mathcal{F}}_{\mathsf{X}}^{k}=\texttt{Seperate}(\hat{\mathcal{S}}_{\mathsf{Concat}}^{k}+\texttt{Inverse}(\hat{\mathcal{S}}_{\mathsf{Inverse}}^{k})).$ (18)
After obtaining the scanned features ¯$\overline{\mathcal{F}}_{\mathsf{rgb}}^{k}$ and ¯$\overline{\mathcal{F}}_{\mathsf{x}}^{k}$, they are multiplied with two scaling parameters and concatenated in the channel dimension. Finally, a linear projection layer is used to reduce the feature shape to $\mathbb{R}^{H_{k}\times W_{k}\times C_{k}}$.
@@ 4.2 Quantitative and Qualitative Results
RGB-D Semantic Segmentation. In Table 2, we compare Sigma against various RGB-D methods to validate its generalization capability. Remarkably, our small model surpasses the performance of CMNeXt [65] while employing only $69.8$M parameters, which is $49.8$M less than CMNeXt. This demonstrates our method’s superior balance between accuracy and efficiency.
@@ 4.3 Ablation Studies
Comparison of Attention Matrix in CroMB. In CroMB, we exchange matrix $C$ of two modalities for cross-modal information interaction. For comparison, we conduct experiments to use matrix $C$, $D$, $B\&C$, and $C\&D$ in CroMB. As shown in Table 4, using the matrix $C$ achieves the optimal performance. This supports our claim that matrix $C$ is used to decode the information from the hidden state and can guide the reconstruction for complementary modality.
@@ Acknowledgement
This work has been funded in part by the Army Research Laboratory (ARL) under grant W911NF-23-2-0007 and DARPA FA8750-23-2-1015. We also appreciate the constructive suggestions provided by Ce Zhang.
@@ Appendix B Daytime and Nighttime Performance
[…]

### v3

@@ 1 Introduction
Previous approaches in multi-modal semantic segmentation rely on Convolutional Neural Networks (CNN) or Vision Transformers (ViT). While CNN-based approaches [12, 24, 67] are known for their scalability and linear complexity, they suffer from the small receptive field limited by the kernel size, leading to the local reductive bias. Besides, CNNs utilize a weight-sharing mechanism across different parts of the input, limiting its flexibility in adapting to unseen or low-quality images. In contrast, ViT-based methods [45, 25, 1] offer enhanced visual modeling by leveraging global receptive fields and dynamic weights. However, their self-attention mechanism leads to quadratic complexity in terms of input sizes [14], raising efficiency concerns. Attempts to improve this efficiency by reducing the dimensions or strides of the processing windows compromise the extent of the receptive fields [59].
@@
Our Sigma integrates a Siamese encoder [55, 56] for modality-specific feature extraction, fusion modules for aggregating information from different modalities, and a decoder for enhancing both spatial and channel-specific information. The encoder backbone utilizes cascaded Visual State Space Blocks (VSSB) to extract multi-scale information from various modalities. Subsequently, the extracted features are inputted to a fusion module at each level, where multi-modal features initially interact through a Cross Mamba Block (CroMB) to augment cross-modal information. Then, the enhanced features are processed by a Concat Mamba Block (ConMB), which employs an attention mechanism to select pertinent information from each modality. Finally, the fused features are sent to multi-level Channel-Aware Visual State Space Blocks (CAVSSB) to effectively capture multi-scale long-range information.
@@ 3.4 Fusion Module
$\hat{\mathcal{F}}_{\mathsf{rgb}}^{k},\;\hat{\mathcal{F}}_{\mathsf{x}}^{k}=\texttt{CroMB}(\mathcal{F}_{\mathsf{rgb}}^{k},\;\mathcal{F}_{\mathsf{x}}^{k}),$ (6)
@@
Here, $\hat{\mathcal{F}}_{\mathsf{rgb}}^{k},\hat{\mathcal{F}}_{\mathsf{x}}^{k},\text{and }\mathcal{F}_{\mathsf{fuse}}^{k}$ remains the original dimension as $\mathbb{R}^{H_{k}\times W_{k}\times C_{k}}$. Details of CroMB and ConMB are as follows.
@@
CroMB. As demonstrated in the upper part of Fig. 4, two input features are first processed by linear layers and depth-wise convolutions, then sent to the Cross Selective Scan (Cross SS) module. According to the selection mechanism of Mamba mentioned in Sec. 3.1, the system matrices B,CB,\;C and $\mathrm{\Delta}$ are generated to enable the context-aware ability of the model. Here, linear projection layers are utilized to generate the matrices. According to Eq. 4, matrix $C$ is used to decode the information from the hidden state $h_{k}$ to obtain the output $y_{k}$. Inspired by the cross-attention mechanism [3], we facilitate information exchange among multiple selective scan modules. In particular, the process can be represented as:
@@
Specifically, the outputs $\hat{\mathcal{F}}_{\mathsf{rgb}}^{k},\hat{\mathcal{F}}_{\mathsf{x}}^{k}\in\mathbb{R}^{H_{k}\times W_{k}\times C_{k}}$ from CroMB are first processed by linear and depth-wise convolution layers, then sent to the Concat Selective Scan (Concat SS) module. Within the Concat SS module, the two features are first flattened and then concatenated in the sequence length dimension. This provides a sequence $\mathcal{S}_{\mathsf{Concat}}^{k}$ of shape $\mathbb{R}^{(2\times H_{k}\times W_{k})\times C_{k}}$. Besides, to comprehensively capture information from two modalities, we inversely scan the concatenated sequence $\mathcal{S}_{\mathsf{Concat}}^{k}$ to obtain an additional sequence $\mathcal{S}_{\mathsf{Inverse}}^{k}\in\mathbb{R}^{(2\times H_{k}\times W_{k})\times C_{k}}$. Afterwards, each sequence is processed to capture long-range dependencies, obtaining $\hat{\mathcal{S}}_{\mathsf{Concat}}^{k}$ and $\hat{\mathcal{S}}_{\mathsf{Inverse}}^{k}$. Then the inversed sequence output is flipped back and added with the processed sequence. The summed sequence is separated to recover two outputs. This process can be represented as:
$\widetilde{\mathcal{F}}_{\mathsf{rgb}}^{k}=\texttt{DWConv}(\texttt{Linear}(\hat{\mathcal{F}}_{\mathsf{rgb}}^{k})),$ (13)
$\widetilde{\mathcal{F}}_{\mathsf{x}}^{k}=\texttt{DWConv}(\texttt{Linear}(\hat{\mathcal{F}}_{\mathsf{x}}^{k})),$ (14)
@@
$\overline{\mathcal{F}}_{\mathsf{RGB}}^{k},\overline{\mathcal{F}}_{\mathsf{X}}^{k}=\texttt{Seperate}(\hat{\mathcal{S}}_{\mathsf{Concat}}^{k}+\texttt{Inverse}(\hat{\mathcal{S}}_{\mathsf{Inverse}}^{k})).$ (18)
After obtaining the scanned features $\overline{\mathcal{F}}_{\mathsf{rgb}}^{k}$ and $\overline{\mathcal{F}}_{\mathsf{x}}^{k}$, they are multiplied with two scaling parameters and concatenated in the channel dimension. Finally, a linear projection layer is used to reduce the feature shape to $\mathbb{R}^{H_{k}\times W_{k}\times C_{k}}$.
@@ 4.2 Quantitative and Qualitative Results
RGB-D Semantic Segmentation. In Table 2, we compare Sigma against various RGB-D methods to validate its generalization capability. Remarkably, our small model surpasses the performance of CMNeXt [68] while employing only 69.869.8M parameters, which is 49.849.8M less than CMNeXt. This demonstrates our method’s superior balance between accuracy and efficiency.
@@ 4.3 Ablation Studies
Comparison of Attention Matrix in CroMB. In CroMB, we exchange matrix $C$ of two modalities for cross-modal information interaction. For comparison, we conduct experiments to use matrix $C$, $D$, B&CB\&C, and C&DC\&D in CroMB. As shown in Table 4, using the matrix $C$ achieves the optimal performance. This supports our claim that matrix $C$ is used to decode the information from the hidden state and can guide the reconstruction for complementary modality.
@@ Acknowledgement
This work has been funded in part by the Army Research Laboratory (ARL) award W911NF-23-2-0007, DARPA award FA8750-23-2-1015, and ONR award N00014-23-1-2840. We also appreciate the constructive suggestions provided by Ce Zhang.
@@ Appendix B Daytime and Nighttime Performance
Table B2: Performance comparison on daytime and nighttime MFNet [19] dataset. We use mIoU ($\%$) for evaluation.